# Sequences that do frame reconstruction

### v2

@@ 5. A Paley–Wiener type stability theorem for bounded sequences that do frame reconstruction
Let $\mathcal{H}$ be a Hilbert space and suppose $\{f_{n}\}_{n\in I}\subseteq\mathcal{H}$ where $I$ is $\mathbb{N}$ or $\mathbb{Z}$ does (unconditional) frame reconstruction with $B$ and
@@
$\sum_{n\in I}\|f_{n}-h_{n}\|<\sqrt{M^{2}+\frac{1}{\|B\|}}-M.$
Then $\{h_{n}\}$ also does (unconditional) frame reconstruction with operator $(T^{-1})^{*}B$, where $T:\mathcal{H}\to\mathcal{H}$
@@
Now let $\lambda=\sqrt{M^{2}+\frac{1}{\|B\|}}-M$. In can be shown in a similar way that for all $f\in\mathcal{H}$,
$\left\|\sum_{n=0}^{\infty}\langle f,Bh_{n}\rangle h_{n}-\sum_{n=0}^{\infty}\langle f,Bf_{n}\rangle f_{n}\right\|<\lambda(2M+\lambda)\|B\|\|f\|.$

### v3

@@ 5. A Paley–Wiener type stability theorem for bounded sequences that do frame reconstruction
Let $\mathcal{H}$ be a non-trivial Hilbert space and suppose $\{f_{n}\}_{n\in I}\subseteq\mathcal{H}$ where $I$ is $\mathbb{N}$ or $\mathbb{Z}$ does (unconditional) frame reconstruction with $B$ and
@@
$\sum_{n\in I}\|f_{n}-h_{n}\|\leq\frac{1}{2}\left(\sqrt{M^{2}+\frac{1}{\|B\|}}-M\right).$
Then $\{h_{n}\}$ also does (unconditional) frame reconstruction with operator
$(T^{-1})^{*}B$, where $T:\mathcal{H}\to\mathcal{H}$
@@
Now let $\lambda=\frac{1}{2}\left(\sqrt{M^{2}+\frac{1}{\|B\|}}-M\right)$. Consider the following for any $f\in\mathcal{H}$,
$\left\|\sum_{n=0}^{\infty}\langle f,Bh_{n}\rangle h_{n}-\sum_{n=0}^{\infty}\langle f,Bf_{n}\rangle f_{n}\right\|\leq$
$\left\|\sum_{n=0}^{\infty}\langle f,Bh_{n}\rangle(h_{n}-f_{n})\right\|+\left\|\sum_{n=0}^{\infty}\langle f,B(h_{n}-f_{n})\rangle f_{n}\right\|\leq$
$\lambda\|f\|\|B\|sup_{n}\|h_{n}\|+\|B\|\|f\|\lambda M\leq\lambda\|f\|\|B\|(M+\lambda)+\|B\|\|f\|\lambda M=$
$\lambda\|B\|((M+\lambda)+M)\|f\|.$